# Matchings and loose cycles in the semirandom hypergraph model

### v1

@@ Abstract.
We study the 2-offer semirandom 3-uniform hypergraph model on $n$ vertices. At each step, we are presented with 2 uniformly random vertices. We choose any other vertex, thus creating a hyperedge of size 3. We show a strategy that constructs a perfect matching, and another that constructs a loose Hamilton cycle, both succeeding asymptotically almost surely within $\Theta(n)$ steps. Both results extend to $s$-uniform hypergraphs. Much of the analysis is done on an auxiliary graph that is a uniform $k$-out subgraph of a random bipartite graph, and this tool may be useful in other contexts.
@@ 1.1. The model
We consider the semirandom graph process suggested by Peleg Michaeli and introduced formally in [3]. This model, that can be viewed as a “one player game”, generalizes many well-known and important processes such as the Erdős-Rényi random graph process and the $k$-out process, and has received a lot of attention recently. We concentrate on its natural generalization to hypergraphs, as proposed in [1].
@@
Our second main result is a somewhat similar strategy to construct a 3-uniform hypergraph $H$ on $n$ vertices, $n$ divisible by 2, so that $H$ will contain contain a loose Hamilton cycle.
@@ 1.2. Outline
The next section gives some background on the semirandom process. Section 4 proves Theorem 1 by presenting a 2-offer strategy for constructing a 3-uniform hypergraph matching, and analysing it. Subsection 4.4 proves the extension to $s$-uniform hypergraphs. This is a simple modification of the argument for the 3-uniform case but requires reference to the strategy: Theorem 2 is not a black-box corollary of Theorem 1. Section 5 proves Theorems 3 and 4. The strategy for constructing a loose Hamilton cycle is different in its details from that for matching, but has the same structure and shares some key elements.
@@ 2. Background
Since the semirandom process is still relatively new, let us highlight a few results on this model, focusing first on perfect matchings.
@@
In the very first paper [3], it was shown that the semirandom process on graphs is general enough to simulate several well-studied random graph models (using suitable strategies), including the extensively studied $k$-out process. In the $k$-out process on $n$ vertices, each vertex independently connects to $k$ randomly selected vertices, resulting in a random graph with $kn$ edges (see, for example, Chapter 18 in [16]). This is easily simulated with the semirandom process by “choosing” vertex 1 to accompany each of the first $k$ offered vertices, choosing vertex 2 for the next $k$, etc.
@@
Since the $2$-out process has a perfect matching a.a.s. [9], we immediately get that a.a.s. one can create a perfect matching in $(2+o(1))n$ rounds. Using the semirandom process to construct a different sort of random graph known to have a perfect matching a.a.s. [17], the upper bound can be improved to $(1+2/e+o(1))n<1.73576n$ [3]. This bound was subsequently improved by investigating a fully adaptive algorithm. The current best upper bound is $1.20524n$ [13]. On the other hand, [3] observed that $(\ln(2)+o(1))n>0.69314n$ is a lower bound on the number of rounds needed to create a perfect matching, and this has since been improved to $0.93261n$ [13].
One approach to constructing perfect matchings in the 2-offer 3-uniform model would be to use a random $k$-out hypergraph, i.e., a hypergraph with $k$ random edges on (and associated with) each vertex. This is easily simulated in the semirandom process just as for graphs, choosing vertex 1 with each of the first $k$ offers, etc. It is natural to expect that such a hypergraph has a perfect matching, but it is only known to have a perfect fractional matching [5]. As said in [5], proving that it has a perfect matching is “likely to be difficult, as [it] would imply the … resolution of Shamir’s Problem, as well as [a] natural guess regarding a beautiful problem of Frieze and Sorkin [7].” Our approach is quite different.
Let us also briefly comment on other results on the semirandom graph model. Unsurprisingly, Hamilton cycles have received a lot of attention [3, 11, 12, 8, 6], and currently the best upper and lower bounds are $1.81701n$ and $1.26575n$ [6]. More general spanning subgraphs were considered in [2]. It showed that any graph with maximum degree $\Delta$ can be constructed a.a.s. in $(3\Delta/2+o(\Delta))n$ rounds; it also showed that if $\Delta=\omega(\log(n))$, this upper bound improves to $(\Delta/2+o(\Delta))n$ rounds, close to the trivial lower bound of $\Delta n/2$. Note, however, that these results are meaningful only for $\Delta\rightarrow\infty$; for any fixed $\Delta$ they say only that $O(n)$ rounds suffice. Cliques, chromatic number, and independent sets were considered in [10].
@@
Finally, while our interest is in hypergraphs, let us mention some other variants of the semirandom process. In [18], sharp thresholds were studied for a more general class of processes that includes the semirandom process. In [4], a random spanning tree of $K_{n}$ is presented, and the player keeps one of the edges. In [14], vertices are presented by the process in a random permutation. In [20], the process presents $k$ random vertices, and to create an edge the player selects one of them, and freely chooses a second vertex.
@@ 3. Outline, and lemmas on an auxiliary bipartite graph
A uniformly random $k$-out bipartite multigraph (or just “uniform $k$-out” graph) on disjoint node sets $\mathcal{A},\mathcal{B}$ is a directed graph $\mathcal{D}$ in which every node in $\mathcal{A}$ has outedges to $k$ nodes in $\mathcal{B}$ chosen uniformly at random with replacement, and symmetrically for $\mathcal{B}$.
@@
Phase 1 will form a uniform $k$-out graph from a subset of a set of uniformly random edges on $\mathcal{A}\times\mathcal{B}$. Because the edges supplied are uniformly random, a few nodes are bound to have too few edges, but Lemma 6 shows that we can get a uniform $k$-out graph $\mathcal{D}$ on most of the nodes in $\mathcal{A}$ and $\mathcal{B}$. The few nodes missing from $\mathcal{D}$ will correspond to points of the hypergraph we will have to deal with specially.
@@
Let $X_{1},\ldots X_{n}$ be independent random variables with $X_{k}$ taking values in a set $A_{k}$ for each $k$. Suppose that the (measurable) function $f\colon\prod A_{k}\rightarrow\mathbb{R}$ satisfies $\left|{f(x)-f^{\prime}(x)}\right|\leq c_{k}$ whenever the vectors $x$ and $x^{\prime}$ differ only in the $k$th coordinate. Let $Y$ be the random variable $f(X_{1},\ldots,X_{n})$. Then for any $t>0$, $\Pr(\left|{Y-\mathbb{E}Y}\right|\geq t)\leq 2\exp(-2t^{2}/\sum c_{k}^{2})$.
@@
Fix $\delta=\min\left\{\epsilon/2,(2e^{2}\binom{k+3}{4})^{-1/2}\right\}$. Let $\lambda$ be the value such that the probability that a Poisson random variable $Z\sim\operatorname{Po}(\lambda)$ is less than $k+3$ is $\delta/4$; let us write this as $\Pr(\operatorname{Po}(\lambda)<k+3)=\delta/4$. Choose $C=2.1\lambda$.
Assume that $\operatorname{Po}(2\lambda N)\leq CN$; the probability that this fails to hold is exponentially small in $N$. Randomly orient the first $\operatorname{Po}(2\lambda N)$ edges of $\Psi$, so there are $\operatorname{Po}(\lambda N)$ edges out of each part.
Suppose part $\mathcal{A}$ has $m$ nodes. Choose $t$ to be the power of 2 for which $\frac{1}{2}N/t<m\leq N/t$. The number of outedges $Z(v)$ from a node $v\in\mathcal{A}$ is Poisson with intensity $\lambda N/m\geq\lambda N/(N/t)=t\lambda$. We may thus write $Z(v)$ as $Z(v)=Z_{1}(v)+\cdots+Z_{t}(v)+Z^{\prime}(v)$ where every $Z_{i}(v)\sim\operatorname{Po}(\lambda)$, $Z^{\prime}(v)$ is also Poisson, and all are independent. The event $Z(v)<k+3$ occurs only if every $Z_{i}(v)<k+3$.
@@
Then, with exponentially small failure probability, all but a $1-\delta$ fraction of vertices in both $\mathcal{A}$ and (by the same argument, and a union bound) in $\mathcal{B}$ have at least $k+3$ outedges. For any node with more than $k+3$ outedges, randomly delete some to leave exactly $k+3$.
Let $\mathcal{D}_{0}$ be the resulting directed bipartite graph. Let $X_{0}$ be the vertices failing to have outdegree $k+3$. Let $\mathcal{D}$ be the $k$-core of $\mathcal{D}_{0}\setminus X_{0}$. $\mathcal{D}$ can be constructed from $\mathcal{D}_{0}$ by initialising $X=X_{0}$, successively adding to $X$ any node with less than $k$ outneighbours in $\mathcal{D}_{0}\setminus X$, or equivalently with at least $4$ outneighbours in $X$. When there are no more such vertices, $\mathcal{D}$ is the subgraph of $\mathcal{D}_{0}$ induced by $\mathcal{A}^{\prime}=\mathcal{A}\setminus X$ and $\mathcal{B}^{\prime}=\mathcal{B}\setminus X$.
We claim that, at the end, $\left|{X\cap\mathcal{A}}\right|<2\delta\left|{\mathcal{A}}\right|$ and $\left|{X\cap\mathcal{B}}\right|<2\delta\left|{\mathcal{B}}\right|$ a.a.s. By choice of $\delta$, this condition immediately implies the lemma’s conclusion. Suppose the condition does not hold. At any time, let $S$ and $T$ be the vertices added in parts $\mathcal{A}$ and, respectively, $\mathcal{B}$ that bring $X_{0}$ to $X$. (I.e., $S=(X\cap\mathcal{A})\setminus(X_{0}\cap\mathcal{A})$.) For the condition to be violated, at the end we must have $\left|{S}\right|\geq\delta\left|{\mathcal{A}}\right|$ or $\left|{T}\right|\geq\delta\left|{\mathcal{B}}\right|$. Consider the first moment that either event first occurs. There are two symmetric cases, and we consider just the one where, at that moment, $\left|{S}\right|=\delta\left|{\mathcal{A}}\right|$ (up to integrality) while $\left|{T}\right|<\delta\left|{\mathcal{B}}\right|$. Note that every node in $S$ has at least 4 outneighbours in $T\cup(X_{0}\cap\mathcal{B})$. This remains true if we expand $T$ arbitrarily (but disjoint from $X_{0}$) so that $\left|{T}\right|=\delta\left|{\mathcal{B}}\right|$. The probability that any such pair $S,T$ exists is at most the expected number of such pairs, over all fixed subsets and subject to the randomness of $\mathcal{D}$. For a fixed $v\in S$, the probability $v$ has at least $4$ outneighbours in $T$ is at most $\binom{k+3}{4}\left({\left|{T}\right|}/{\left|{\mathcal{B}}\right|}\right)^{4}=\binom{k+3}{4}\delta^{4}$. Then, the expected number of pairs $S,T$ is at most
$\binom{N}{\delta N}\binom{N}{\delta N}\left(\binom{k+3}{4}\delta^{4}\right)^{\delta N}$ $\leq\left(\frac{eN}{\delta N}\right)^{2\cdot\delta N}\left(\binom{k+3}{4}\delta^{4}\right)^{\delta N}$
$=\left(e^{2}\binom{k+3}{4}\delta^{2}\right)^{\delta N}\leq(1/2)^{\delta N}=o(1),$
since we chose $\delta$ to ensure that $e^{2}\binom{k+3}{4}\delta^{2}\leq 1/2$.
@@
For any constants $k$ and $d$, in a uniformly random $k$-out digraph with parts $\mathcal{A}$ and $\mathcal{B}$ both of size $\Theta(N)$, the probability that $d$ random nodes in $\mathcal{B}$ have a common inneighbour in $\mathcal{A}$ is $O(1/N)$, and the symmetric condition also holds.
@@
If $d$ nodes in $\mathcal{B}$ have $a\in\mathcal{A}$ as a common inneighbour, then at least two of them must be in $N^{+}(a)$. Since $\left|{N^{+}(a)}\right|=k$ while $\left|{\mathcal{B}}\right|=\Theta(N)$, this has probability $O(1/N^{2})$. Taking the union bound over all $a\in\mathcal{A}$, the probability the $d$ nodes have any common inneighbour is $O(1/N)$. ∎
@@
Note:
@@
  • The lemma implies that sets of size $\delta N$ expand by a factor at most $\eta_{c,k}(\delta)/\delta$. This is false for smaller sets: there will typically be a handful of vertices of arbitrarily high indegree.
  • The proof below uses the first-moment method. A slightly smaller $\eta_{c,k}(\delta)$ can be obtained by choosing (in the Poisson limit) a degree threshold above which there are at least $\delta N$ vertices (thus encompassing the worst-case set $S$), and considering the total of those vertices’ degrees (concentrated, and at least $\left|{N^{-}(S)}\right|$).

### v2

# Perfect matchings and loose Hamilton cycles in the semirandom hypergraph model

@@ Abstract.
We study the 2-offer semirandom 3-uniform hypergraph model on $n$ vertices. At each step, we are presented with 2 uniformly random vertices. We choose any other vertex, thus creating a hyperedge of size 3. We show a strategy that constructs a perfect matching, and another that constructs a loose Hamilton cycle, both succeeding asymptotically almost surely within $\Theta(n)$ steps. Both results extend to $s$-uniform hypergraphs. The challenges with hypergraphs, and our methods, are qualitatively different from what has been seen for semirandom graphs. Much of our analysis is done on an auxiliary graph that is a uniform $k$-out subgraph of a random bipartite graph, and this tool may be useful in other contexts.
@@ 1.1. The model
The canonical Erdős-Rényi random graph model [8] has been studied exhaustively since the 1960s. It is deeply understood, and by now a great deal of attention has been devoted to variants and special cases, including hypergraphs, sparse graphs, planar graphs, and maker-breaker games. A version of the Erdős-Rényi model is that, given $n$ vertices, edges arrive, one per round, uniformly at random. One variant, suggested around 2000 by Achlioptas and first analyzed in [5], is that edges arrive two at a time, and just one of the two is added to the graph.
The semirandom graph process we consider here has a similar character. Instead of being offered a pair of edges, we are offered a vertex, and we may add any edge on that vertex; another way to view it is that we are offered a star, and choose one of its edges to add. This process was suggested by Peleg Michaeli, introduced formally in [4] in 2020, and has already received a good deal of attention. It can be viewed as a “one player game”, and generalizes several well-known and important processes including the Erdős-Rényi random graph process itself, and the “$k$-out process” (about which we will say more). We concentrate on its natural generalization to hypergraphs, as proposed in [1].
@@
1.2. Our contribution
Our contribution is to resolve some natural first questions for the hypergraph semirandom process, notably in the case where each offer specifies two vertices rather than just one. We believe that this is the first work on semirandom hypergraphs that is qualitatively different from anything that has been done for semirandom graphs. In this work, the property $\mathcal{P}$ is that the semirandom hypergraph should contain as a subgraph a desired structure, namely a perfect matching or a loose Hamilton cycle. Constructing such spanning structures has also been one focus of the study of semirandom graphs.
A main point of difference is that for graphs, it has always been possible to quickly and simply construct most of the structure desired, then cleverly fill in the final gaps. That approach does not seem to work here. Instead, we first generate a large absorbing structure of some sort, build a small part of the desired structure using the few vertices absent from the absorber, then use the absorber to complete the desired structure. Also, while the semirandom work on graphs has been able to address the desired graph directly, we found it necessary to work in an auxiliary graph.
To suggest the profoundly greater challenges of semirandom hypergraphs compared with graphs, let us mention an approach to matching that works for graphs but seemingly not for hypergraphs. It is easy to semirandomly construct a random $k$-out graph or hypergraph; we discuss this further in Section 2. In the graph case, by [12], the $k$-out almost surely contains a perfect matching. In the hypergraph case, by contrast, Devlin and Kahn [7] observe that the corresponding statement is “almost certainly correct [but] likely to be difficult [to show]”, as it is stronger than the 2008 resolution of Shamir’s Problem by [19] and would imply a (still open) “natural guess regarding a beautiful problem of Frieze and Sorkin” (from [10]). Devlin and Kahn make progress towards this by proving the presence of a perfect fractional matching, but that is not sufficient for our purpose. Even if this approach did work, it would likely be existential; we prefer (and obtain) an efficient construction.
To avoid confusion, we remark that our successful approach to hypergraph matching has some superficial similarity to this. We are able to construct a $k$-out auxiliary graph (not hypergraph) on most (unfortunately not all) of the vertices to give an absorbing structure that guarantees a hypergraph matching after the exceptional vertices are dealt with.
1.3. Our results
@@
Our second main result is a somewhat similar strategy to construct a 3-uniform hypergraph $H$ on $n$ vertices, $n$ divisible by 2, so that $H$ will contain a loose Hamilton cycle.
@@
For both perfect matchings and Hamilton cycles, the 1-offer results follow as corollaries of the 2-offer ones but, unsurprisingly, can be obtained much more easily. One way to do so is touched upon in a footnote near the start of Section 3: where for both matchings and cycles our 2-offer constructions use a three-phase strategy, for the 1-offer model, a simple version of the first phase suffices. These 1-offer constructions are presented as warm-ups near the start of Section 4 (addressing matchings) and Section 5 (Hamilton cycles).
@@ 1.4. Outline
Section 2 gives some background on the semirandom process. Section 3 sketches the approach we will take. The approach relies greatly on a certain auxiliary graph, and the section motivates, states, and proves the lemmas governing the auxiliary graph. (Roughly: a random bipartite graph contains a large uniformly random $k$-out subgraph. A.a.s., every large subgraph of such a graph, if it has minimum degree at least $k-1$ and parts of equal size, contains a perfect matching.) Section 4 proves Theorem 1 by presenting a 2-offer strategy for constructing a 3-uniform hypergraph matching, and analysing it. Subsection 4.4 proves the extension to $s$-uniform hypergraphs. This is a simple modification of the argument for the 3-uniform case but requires reference to the strategy: Theorem 2 is not a black-box corollary of Theorem 1. Section 5 proves Theorems 3 and 4. The strategy for constructing a loose Hamilton cycle is different in its details from that for matching, with some new complications, but it has the same general structure and shares many key elements.
@@ 2. Background
We now describe some of the main results in semirandom graphs, especially in how they relate to the present work.
A prominent result on semirandom graphs concerns constructing a graph $G$ to contain a given spanning structure $H$, such as a perfect matching or Hamilton cycle [3]. It shows that for a spanner $H$ with maximum degree $\Delta$, this can be done in time $(3\Delta/2+o(\Delta))n$. Roughly speaking, the construction works by fixing an arbitrary correspondence between the vertices of $G$ and $H$, trying to build in $G$ the edges between each vertex $v$ and its desired neighbours $N_{H}(v)$, and (cleverly) patching up the vertices for which this fails. For the 2-offer 3-uniform semirandom model, this approach fails as soon as it fixes a correspondence and thus dictates the edges desired: to construct a given edge $(u,v,w)$ requires an offer of some two of these vertices, and in time $O(n)$ the chance of this occurring is only $O(1/n)$.
Even for graphs, a limitation of [3] is that the $o(\Delta)$ term means that the coefficient of the linear running time is unknown for any fixed $\Delta$. It therefore does not resolve efficient constructions for low-degree graphs of common interest, prominently including perfect matchings and Hamilton cycles.
@@
In the paper introducing the semirandom model, [4], it was shown that the semirandom process on graphs is general enough to simulate several well-studied random graph models (using suitable strategies), including the extensively studied $k$-out process. Since the $k$-out model is essential to the work here, let us elaborate.
In the $k$-out process on $n$ vertices, each vertex independently connects to $k$ randomly selected vertices, resulting in a random graph with $kn$ edges (see, for example, [20, Chapter 18]). To capture the association between vertices and edges, it is often convenient to think of the process’s result as a $k$-out random digraph with $k$ random directed edges out of each vertex. This is easily simulated with the semirandom process by “choosing” vertex 1 to accompany each of the first $k$ offered vertices (constructing a directed edge from vertex 1 to each of them), choosing vertex 2 for the next $k$, etc. (Details, such as the possibility of being offered vertex 1 when trying to construct edges out of vertex 1, are easily resolved and uninteresting.)
@@
A linear-time semirandom construction of a perfect matching follows from this, since a.a.s. the $2$-out process has a perfect matching [12]. Using the semirandom process to construct a different sort of random graph known to have a perfect matching a.a.s. (see [21]), the upper bound can be improved to $(1+2/e+o(1))n<1.73576n$ [4].
In the 2-offer 3-uniform semirandom model it is equally easy to generate a random $k$-out hypergraph, i.e., a hypergraph with $k$ random edges on (and associated with) each vertex. Simply, to each of the first $k$ offered pairs add vertex 1, and so on. However, as discussed in the introduction (Section 1.2), it is not known whether a random $k$-out hypergraph a.a.s. contains a perfect matching, and our approach to constructing a matching (and loose Hamilton cycle) is entirely different.
For both perfect matchings and Hamilton cycles, researchers have sought the most efficient semirandom constructions possible. For matchings, currently the most efficient construction uses time $1.20524n$ [16]. On the other hand, [4] observed that $(\ln(2)+o(1))n>0.69314n$ is a lower bound on the number of rounds needed to create a perfect matching (indeed, to give each vertex nonzero degree), and this has since been improved to $0.93261n$ [16].
Hamilton cycles in semirandom graphs have been studied in [4, 14, 15, 11, 9], and currently the best upper and lower bounds are $1.81701n$ and $1.26575n$ [9].
Another line of research on the semirandom model is in constructing a graph containing a given, fixed graph $H$. Reminiscent of the derivation of thresholds for the presence of a fixed subgraph in a (usual) random graph, the task is quite different from the one studied here. For one thing, it can be accomplished in sublinear time: where $H$ has degeneracy $d$, a copy can be constructed in time just larger than $n^{(d-1)/d}$ [4], but cannot be constructed in time just smaller than that [1]. The construction method is different from what has been done for spanning subgraphs, and what is done here: a copy of $H$ is constructed vertex by vertex, in an order in which each vertex $v$ needs to connect to at most $d$ previous ones, using all offered vertices to try to fulfill the role of $v$. The lower bound in [1] is more complicated but uses the same general ideas. We mention this work primarily because [1] extends these graph results to hypergraphs, providing one of the few results for semirandom hypergraphs. However, the extension to the 1-offer model is straightforward, while for 2 or more offers the understanding is incomplete, with the upper and lower bounds matching only for some special cases [2].
@@
We briefly mention a few other results. Semirandom constructions to achieve the largest possible clique, largest possible chromatic number, and smallest possible independent set are studied in [13]. Sharp thresholds for the semirandom process and a more general class of processes were studied in [22]. Variants of the semirandom process have also been explored. In [17], the first $n$ vertices offered are all distinct, as given by a random permutation; the next $n$ vertices are given by another permutation; and so on. In [6], a random spanning tree of $K_{n}$ is presented, and the player keeps one of the edges. In [24], $k$ random vertices are offered, and the player selects one of them, and freely chooses a second vertex.
@@ 3. Outline, and lemmas on an auxiliary bipartite graph
A uniformly random $k$-out bipartite multigraph (or just “uniform $k$-out” graph) on disjoint node sets $\mathcal{A},\mathcal{B}$ is a directed graph $\mathcal{D}$ in which every node in $\mathcal{A}$ has outedges to $k$ nodes in $\mathcal{B}$ chosen uniformly at random with replacement, and symmetrically for $\mathcal{B}$. The edge directions are generally ignored, here in particular, except as a convenient way to associate a vertex with “its” $k$ edges (namely, its $k$ out-edges).
@@
Phase 1 will form a uniform $k$-out auxiliary graph from a subset of a set of uniformly random edges on $\mathcal{A}\times\mathcal{B}$. Because the edges supplied are uniformly random, a few nodes are bound to have too few edges, but Lemma 6 shows that we can get a uniform $k$-out graph $\mathcal{D}$ on most of the nodes in $\mathcal{A}$ and $\mathcal{B}$. The few nodes missing from $\mathcal{D}$ will correspond to points of the hypergraph we will have to deal with specially. [^1] In the 1-offer model we could easily build a $k$-out graph on all the nodes, whereupon for $k\geq 2$ the auxiliary graph a.a.s. has a perfect matching which, in turn, corresponds — according to the case — to a hypergraph perfect matching, or a hypergraph loose Hamilton cycle. We will note these cases as we come to them in early in Section 4 and 5.
@@
Let $X_{1},\ldots X_{n}$ be independent random variables with $X_{k}$ taking values in a set $A_{k}$ for each $k$. Suppose that the (measurable) function $f\colon\prod A_{k}\rightarrow\mathbb{R}$ satisfies $\left|{f(x)-f(x^{\prime})}\right|\leq c_{k}$ whenever the vectors $x$ and $x^{\prime}$ differ only in the $k$th coordinate. Let $Y$ be the random variable $f(X_{1},\ldots,X_{n})$. Then for any $t>0$, $\Pr(\left|{Y-\mathbb{E}Y}\right|\geq t)\leq 2\exp(-2t^{2}/\sum c_{k}^{2})$.
@@
The directedness of this graph is, as touched in on Section 2, just a convenient way to express that each node has exactly $k$ out-edges, whose opposite endpoints are independent.
@@
We use a Poissonization argument. Fix $\delta=\min\left\{\epsilon/2,0.99\,(16e^{2}\binom{k+3}{4})^{-1/2}\right\}$. Let $\lambda$ be the value such that the probability that a Poisson random variable $Z\sim\operatorname{Po}(\lambda)$ is less than $k+3$ is $\delta/4$; let us write this as $\Pr(\operatorname{Po}(\lambda)<k+3)=\delta/4$. Choose $C=2.1\lambda$.
Assume that $\operatorname{Po}(2\lambda N)\leq CN$; the probability that this fails to hold is exponentially small in $N$. Attend only to the first $\operatorname{Po}(2\lambda N)$ edges of $\Psi$, ignoring the later ones.
Randomly orient these edges. By the splitting property of the Poisson distribution, the numbers of edges out of parts $\mathcal{A}$ and $\mathcal{B}$ are both $\operatorname{Po}(\lambda N)$ and they are independent.
Suppose part $\mathcal{A}$ has $m$ nodes. Further split the edges out of $\mathcal{A}$ according to which node of $\mathcal{A}$ they are directed out of. Since the edges are by hypothesis random, by the splitting property of the Poisson, the number of edges $Z(v)$ out of each $v\in\mathcal{A}$ is distributed as $Z(v)\sim\operatorname{Po}(\lambda N/m)$, and the $Z(v)$ are mutually independent over all $v\in\mathcal{A}$.
Choose $t$ to be the power of 2 for which $\frac{1}{2}N/t<m\leq N/t$. Splitting the Poisson again, write $Z(v)$ as $Z(v)=Z_{1}(v)+\cdots+Z_{t}(v)+Z^{\prime}(v)$ where every $Z_{i}(v)\sim\operatorname{Po}(\lambda)$, $Z^{\prime}(v)$ is also Poisson, and all are independent. The event $Z(v)<k+3$ occurs only if every $Z_{i}(v)<k+3$.
@@
Then, with exponentially small failure probability, all but a $1-\delta$ fraction of nodes in both $\mathcal{A}$ and (by the same argument, and a union bound) in $\mathcal{B}$ have at least $k+3$ outedges. For any node with more than $k+3$ outedges, randomly delete some to leave exactly $k+3$.
Let $\mathcal{D}_{0}$ be the resulting directed bipartite graph. Let $X_{0}$ be the nodes failing to have outdegree $k+3$. Let $\mathcal{D}$ be the $k$-core of $\mathcal{D}_{0}\setminus X_{0}$. $\mathcal{D}$ can be constructed from $\mathcal{D}_{0}$ by initialising $X=X_{0}$, successively adding to $X$ any node with less than $k$ outneighbours in $\mathcal{D}_{0}\setminus X$, or equivalently with at least $4$ outneighbours in $X$. When there are no more such nodes, $\mathcal{D}$ is the subgraph of $\mathcal{D}_{0}$ induced by $\mathcal{A}^{\prime}=\mathcal{A}\setminus X$ and $\mathcal{B}^{\prime}=\mathcal{B}\setminus X$.
We claim that, at the end, $\left|{X\cap\mathcal{A}}\right|<2\delta\left|{\mathcal{A}}\right|$ and $\left|{X\cap\mathcal{B}}\right|<2\delta\left|{\mathcal{B}}\right|$ a.a.s. By hypothesis, $\delta\leq\epsilon/2$, so this condition immediately implies the lemma’s conclusion. Suppose the condition does not hold. At any time, let $S$ and $T$ be the nodes added in parts $\mathcal{A}$ and, respectively, $\mathcal{B}$ that bring $X_{0}$ to $X$. (I.e., $S=(X\cap\mathcal{A})\setminus(X_{0}\cap\mathcal{A})$.) For the condition to be violated, at the end we must have $\left|{S}\right|\geq\delta\left|{\mathcal{A}}\right|$ or $\left|{T}\right|\geq\delta\left|{\mathcal{B}}\right|$. Consider the first moment that either event first occurs. There are two symmetric cases, and we consider just the one where, at that moment, $\left|{S}\right|=\delta\left|{\mathcal{A}}\right|$ (up to integrality) while $\left|{T}\right|<\delta\left|{\mathcal{B}}\right|$. Note that every node in $S$ has at least 4 outneighbours in $T\cup(X_{0}\cap\mathcal{B})$. This remains true if we expand $T$ arbitrarily (but disjoint from $X_{0}$) so that $\left|{T}\right|=\delta\left|{\mathcal{B}}\right|$. The probability that any such pair $S,T$ exists is at most the expected number of such pairs, over all fixed subsets and subject to the randomness of $\mathcal{D}$. For a fixed $v\in S$, the probability $v$ has at least $4$ outneighbours in $T\cup(X_{0}\cap\mathcal{B})$ is at most $\binom{k+3}{4}\left({\left|{T\cup(X_{0}\cap\mathcal{B})}\right|\;}/{\;\left|{\mathcal{B}}\right|}\right)^{4}=\binom{k+3}{4}(2\delta)^{4}$. Then, the expected number of pairs $S,T$ is at most
$\binom{N}{\delta N}\binom{N}{\delta N}\left(\binom{k+3}{4}(2\delta)^{4}\right)^{\delta N}$ $\leq\left(\frac{eN}{\delta N}\right)^{2\cdot\delta N}\left(\binom{k+3}{4}(2\delta)^{4}\right)^{\delta N}$
$=\left(e^{2}\binom{k+3}{4}2^{4}\delta^{2}\right)^{\delta N}\leq 0.99^{\delta N}=o(1),$
using the hypothesis on $\delta$.
@@
For any constants $k$ and $d$, in any $k$-out digraph with parts $\mathcal{A}$ and $\mathcal{B}$ both of size $\Theta(N)$, $d$ random nodes in $\mathcal{B}$ have disjoint inneighbourhoods with probability $1-O(1/N)$, and the same holds for $d$ random nodes in $\mathcal{A}$.
@@
For a given $a\in\mathcal{A}$, if two of the $d$ nodes in $\mathcal{B}$ have $a$ as a common inneighbour, both are in $N^{+}(a)$. Since $\left|{N^{+}(a)}\right|=k$ while $\left|{\mathcal{B}}\right|=\Theta(N)$, this has probability $O(1/N^{2})$. Taking the union bound over all $a\in\mathcal{A}$, the probability that any two $d$ nodes have a common inneighbour is $O(1/N)$. ∎
@@
That is, any set of up to a $\delta$ fraction of the nodes expands to at most an $\eta_{c,k}(\delta)$ fraction. Note:
@@
  • The lemma implies that sets of size $\delta N$ expand by a factor at most $\eta_{c,k}(\delta)/\delta$. This is false for smaller sets: there will typically be a handful of nodes of arbitrarily high indegree.
  • The proof below uses the first-moment method. A slightly smaller $\eta_{c,k}(\delta)$ can be obtained by choosing (in the Poisson limit) a degree threshold above which there are at least $\delta N$ nodes (thus encompassing the worst-case set $S$), and considering the total of those nodes’ degrees (concentrated, and at least $\left|{N^{-}(S)}\right|$).